OpenStudy (anonymous):

If y=61 when x=-305 what is y when x=-315?

OpenStudy (goldphenoix):

You can set up a proportion. \[\frac{ 61 }{ -305 } = \frac{ y }{ -315 }\] Cross Multiply: -305y = 61 x -315 Simplify: -305y = -19215 Divide to solve for Y: Y= 63.

OpenStudy (anonymous):

Thank You